OpenStudy (anonymous):

what is -1 2/3 divided by 5/6 A -1 7/18 B -2 C 2 D -1 1/6

OpenStudy (liv1234):

-1 2/3 = -1/3 -1/3 divided by 5/6 = -1/3 * 6/5

OpenStudy (anonymous):

ok

OpenStudy (liv1234):